OpenStudy (anonymous):

in a bag of small balls 1/4 are green 1/8 are blue 1/12 are yellow and the remaining 26 white how many balls are blu

OpenStudy (anonymous):

Let us first find the fraction of green, blue and yellow balls

OpenStudy (anonymous):

1/4 + 1/8 + 1/12 = 6 / 24 + 3 / 24 + 2 / 24 , common denominator = 11 / 24 , add numerators

OpenStudy (anonymous):

The fraction of white balls is given by 24 / 24 - 11 / 24 = 13 / 24 So the fraction 13 / 24 corresponds to 26 balls. If x is the total number of balls then (13 / 24) of x = 26 balls or (13 / 24) × x = 26 x = 26 × (24 / 13) = 48 , total number of balls

OpenStudy (anonymous):

The fraction of blue balls is 1 / 8 of x. The number of blue balls is given by (1 / 8) of 48 = 6 balls. =)

OpenStudy (anonymous):

yup 6 balls , i just got the answer too

OpenStudy (anonymous):

lol =)